# Ankle Exoskeletons May Hinder Standing Balance in Simple Models of Older and Younger Adults

### v2

@@ 5.1 Stabilizable regions for age and sex adjusted models
Table 2 summarizes the results for the YF, YM, OM, OF, and wOF models with the gravity compensation (GC) and PD controller strategies. The changes in area for the forward and backward velocity regions are listed separately. In the forward velocity region the GC and PD controller slightly reduce the area of the SR in the young adult models but increase the total area in the OF and wOF model. The backward velocity region shows a different trend — the exoskeleton assistance reduces the area of the stabilizable region for all models. The additional damping provided by the PD controller does not significantly mitigate reductions in stabilizable region area. The stabilizable regions for the young male model and the weak older female model are shown in the top two panels of Figure 4, with and without the exoskeleton added.
$v^{+}$ Gravity compensation Proportional-derivative
YM $-0.43$ $-0.37$
YF $-3.14$ $-2.73$
OM $+3.71$ $+4.64$
OF $+8.43$ $+9.56$
wOF $+15.9$ $+17.5$
$v^{-}$ Gravity compensation Proportional-derivative
YM $-16.9$ $-16.5$
YF $-32.7$ $-32.4$
OM $-15.39$ $-14.94$
OF $-19.56$ $-18.92$
wOF $-10.5$ $-10.16$
Table 2: Effect of exoskeleton assistance on total area of stabilizable region for sex and age adjusted models.
5.2 Effect of reduced MT and MRTD
The effect of reducing MT and MRTD in $10\%$ increments can been seen in Figure 5 as percent reduction in area relative to the baseline. Figure 4 shows the YF stabilizable region when MT or MRTD are reduced by $30\%$ and $50\%$ with and without exoskeleton assistance. At positive velocities, reducing maximum torque causes large reductions in SR area but the addition of exoskeleton assistance mitigates the overall effect. For example, a $30\%$ MT reduction results in an SR area that is $20\%$ smaller than the baseline region while there is only a $10\%$ reduction when ankle exoskeleton assistance is included. At negative velocities, the shift in the constraints on human torque arising from (4) overpowers any mitigating effect that the exoskeleton assistance my have on reduced MT.
When MRTD is reduced the effect on the overall SR only becomes noticeable after large reductions of $40\%$ or more at positive velocities. When an exoskeleton is added, however, the effect is amplified. Without the exoskeleton a $30\%$ reduction in MRTD causes a $2.6\%$ reduction in SR area. With an exoskeleton the SR area is reduced by $10.4\%$. At negative velocities, the change in total area is overwhelmingly determined by the change in constraint bounds, as with reduced MT.
Figure 5: Trends in stabilizable region area when MT (circular markers) and MRTD (diamond markers) are independently reduced, with (purple) and without the exoskeleton (orange).
6 Discussion
Figure 6: Boundaries of the SR computed for a simple pendulum using work-energy (W-E) analysis. The regions computed with our method are overlaid in light purple and orange.
The stabilizable regions computed without ankle exoskeleton assistance align with what is known from prior experiments in the literature. Reduced MT has a far stronger effect than MRTD, with even a severe reduction in MRTD contributing to only a $10\%$ reduction in total area. Interestingly, the stabilizable regions computed with ankle exoskeleton assistance suggest that the device may exacerbate MRTD deficits. Reductions in MRTD lead to larger reductions in the total stabilizable region area computed with exoskeleton than without. The same effect is not seen for reduced MT, where the ankle exoskeleton seems to mitigate the reduction in area.
However, as can seen clearly in the Figures 4 and 5, the addition of exoskeleton torque does not uniformly increase or decrease the area of the stabilizable region (SR). We see the following effects in the positive (posterior-anterior) velocity region:
  • The exo increases the SR along the entire backward fall boundary.
  • The exo increases the SR along the forward fall boundary at low velocities.
@@ 6 Discussion
At negative (anterior-posterior) velocities we see similar effects:
  • The exo increases the SR along the forward fall boundary up to nondimensional velocities of approximately $-0.7$.
  • The exo increases the SR boundary along the backward fall boundary at low velocities.
@@
To understand why this may be happening, we will analyze the dynamics of an unconstrained and undamped 2d pendulum, with and without the exoskeleton, e.g. the system with dynamics $\tau_{h}+\tau_{\text{exo}}=ml^{2}\ddot{\theta}+mgl\cos(\theta)$.
We first note that gravity exerts a negative torque when the CoM is behind the ankle and a positive torque when the CoM is ahead of the ankle. We can draw a vertical line at $y=\theta_{\text{ankle}}$ that separates the 2d state space into positive and negative gravity torque regions. This is the solid black line shown in Figure 6.
We also need to define a zero human-torque line.
Definition 5.
Consider first, the no exoskeleton case — The only static standing position where the human does not need to produce a gravity resisting torque is directly above the ankle. Thus, there is a unique ZTL along which total mechanical energy is a constant, and is equal to the potential energy when the center of mass is directly above the ankle. If the initial energy is $E_{i}=mgl\sin{\theta}+\frac{1}{2}ml^{2}\dot{\theta}^{2}$ and the final energy is $E_{f}=mgl\sin{\theta_{\text{ankle}}}$ ) we can compute the zero torque line by letting $E_{i}=E_{f}$, and thus derive an equation for the system velocity as a function of position:
[…]
This is shown as a dashed black line in figure 6.
The ZTL is useful, as it delineates the direction of the optimal human torque. For a trajectory starting below the ZTL, a positive (dorsiflexion) torque must eventually be applied to avoid a backward fall past the heel. For trajectories starting above the zero-torque line, a negative (plantar flexion) torque is eventually necessary to brake quickly enough to prevent the CoM from overshooting past the toe. The boundary of the stabilizable region that is below the ZTL corresponds to the lower bound before a backward fall occurs, whereas the stabilizable region boundary above the ZTL corresponds to forward falls [20].
It is important to note the overlap between the region where dorsiflexion and plantar flexion are required according to the zero torque trajectory, and the direction of the gravitational torque. In large portions of the state space, gravity provides an assistive torque in the optimal direction.
Now let us consider how the addition of an exoskeleton changes the ZTL. For the analysis in this section, we will consider only the case of the gravity compensation exoskeleton. Equation (18) is no longer valid once a gravity compensation torque is added, as the exoskeleton will always apply a torque. To stay on the original ZTL, i.e. the ZTL without an exoskeleton, the human would have to produce a torque resisting the exoskeleton.
Recall that the exoskeleton torque saturates at a negative and positive value. In the case of the gravity compensation controller we can easily calculate the angles at which the postive and negative saturations occur. We can therefore divide the state space into three regions: with positive saturation, where $\tau_{\text{exo}}=\tau_{\text{sat}}$; no saturation, where $\tau_{\text{exo}}=mgl\cos{\theta}$ ); and negative saturation,where $\tau_{\text{exo}}=-\tau_{\text{sat}}$. The ZTL must be computed for each of these regions in a piecewise manner.
is the ZTL in the region of positive exoskeleton saturation. The full ZTL with exoskeleton assistance is shown in Figure 6 as a dashed red line.
Examining the ZTL with the exoskeleton actuation we can see that in the forward velocity half of the phase plane the line is shifted to the left, such that a plantar flexion braking torque is always required to prevent forward falls. Interestingly, the addition of the exoskeleton effectively eliminates the possibility of a backward fall occurring, as long as there is an initial positive velocity and the contact constraints are satisfied. In the negative velocity half of the plane the ZTL is shifted to the right, increasing the area where a dorsiflexion braking torque is needed to prevent a backward fall. Recalling that the maximum dorsiflexion torque magnitude can be as low as $25\%$ of maximum plantar flexion torque it is not surprising that the increased reliance on dorsiflexion also causes the constraint failure edge to shift inward. These shifts in the ZTL mean that the exoskeleton qualitatively changes the optimal human response required to stabilize to quiet standing in large regions of the state space.
@@
As noted earlier, gravity sometimes provides an assistive torque in the no exoskeleton case. Therefore, if the exoskeleton controller is designed to cancel the gravitational torque, we might expect that exoskeleton “assistance” does not uniformly improve stability. On the other hand, the exoskeleton increases the torque bounds of the human user. This is particularly notable in weaker models as foot-ground interaction constraints are the main limiting factor for stronger models. In particular, static stability is increased in weaker models, resulting in a larger functional base of support.
There is therefore a trade-off which we can think of in terms of work. A weak adult wearing the exoskeleton has the ability to do more work, but they must compensate at times for the gravitational torque being canceled out. We can use the work-energy principle to delineate how this tradeoff affects stabilizability.
The net work of the system is the sum of the work done by the human, exoskeleton, and gravity. Note that the net work done by gravity and the exoskeleton will always have opposite signs due to the controller formulation. For most initial conditions reaching the target set requires a reduction in kinetic energy, i.e. negative work,
[…]
By selecting an appropriate final position $\theta_{f}$, and by assuming that the human produces a constant optimal torque, upper and lower boundaries on the stabilizable region can be computed in a piecewise manner on the position-velocity plane. For example, for $\theta_{i}<\theta_{\text{sat}}^{\text{lb}}$ the upper boundary is expressed by the relation
$\tau_{\text{human}}^{\text{max pf}}\cdot(\theta_{\text{sat}}^{\text{lb}}-\theta)+(-mgl)(\sin(\theta_{\text{sat}}^{\text{ub}})-\sin(\theta_{\text{toeEq}}))+$
$mgl\cos(\theta_{\text{toeEq}})(\theta_{\text{sat}}^{\text{ub}}-\theta_{\text{toeEq}}).$
The bounds derived from the work-energy analysis are shown for the system with (purple) and without (orange) the gravity compensating exoskeleton in Figure 6. As can be seen in the figure, the computed BRS for the exo and non-exo systems are well-approximated by these bounds, demonstrating that the changes in the stabilizable region boundaries can be understood from a work-energy perspective. This analysis, however, assumes conservation of energy and ignores contact and input constraints, leading to an overestimation of the true boundaries. Additionally, it cannot generalize to systems with dissipation — such as exoskeleton controllers with damping.
The stabilizable regions computed with our reachability analysis can therefore increase within these work-energy bounds, subject to foot-ground contact and MRTD constraints. The lack of any significant reduction in area in the full-strength YM model is the result of having a large enough range of positive and negative maximum torque and a large enough MRTD such that the exoskeleton’s undesirable behavior can be compensated for quickly. An exoskeleton that can completely compensate for missing MT and MRTD and uses proportional myo-electric control could lead to similar outcomes.
[…]
7 Derivation of Equations of Motion and Constraints
Here we derive the equations of motion and state constraints presented in Section III.
7.1 Equations of Motion
Figure 7: Free body diagram of system
A free body diagram of the system is shown in Figure 7. To derive the constraints that prevent the foot from slipping or tipping, we must first understand what torques and forces are acting on the foot.
In the following, we will denote $\cos{(\theta)}$ ) as $c_{\theta}$ and $\sin{(\theta)}$ ) as $s_{\theta}$ for particularly unwieldy equations. Let $\tau=\tau_{\text{human}}+\tau_{\text{exoskeleton}}$ be the total internal torque at the ankle. Summing the torques for the body segment at the ankle, we have
$\tau=ml_{b}^{2}\ddot{\theta}+mgl_{b}\cos{\theta},$ (22)
where $ml_{b}^{2}$ is the moment of inertia of the body segment. The rotational motion of the body segment can be decomposed into a normal acceleration along the leg
$a^{n}=-l_{b}\dot{\theta}^{2}$ (23)
and a tangential acceleration
$a^{t}=l_{b}\ddot{\theta}.$ (24)
Rotating these accelerations to the x-y coordinate system, we have -
$a^{x}=a^{n}c_{\theta}-a^{t}s_{\theta}=-l_{b}(\dot{\theta}^{2}c_{\theta}+\ddot{\theta}s_{\theta})$ (25)
$a^{y}=a^{n}s_{\theta}+a^{t}c_{\theta}=l_{b}(-\dot{\theta}^{2}s_{\theta}+\ddot{\theta}c_{\theta})$ (26)
These accelerations are caused by a combination of the weight of the body and the contact forces at the ankle joint. Thus we can write the force balance at the ankle as:
$ma^{x}=F_{\text{joint}}^{x}=-ml_{b}(\dot{\theta}^{2}c_{\theta}+\ddot{\theta}s_{\theta})$ (27)
$ma^{y}=F_{\text{joint}}^{y}-W_{\text{body}}=F_{\text{joint}}^{y}-mg=ml_{b}(-\dot{\theta}^{2}s_{\theta}+\ddot{\theta}c_{\theta})$ (28)
Similarly, for the foot segment we can write the force balance as:
$m_{f}a_{f}^{x}=F_{\text{friction}}-F_{\text{joint}}^{x}$ (29)
$m_{f}a_{f}^{y}=F_{\text{normal}}-F_{\text{joint}}^{y}-W_{\text{foot}}=F_{\text{normal}}-F_{\text{joint}}^{y}-m_{f}g$ (30)
where $m_{f}$ is the mass of the foot. When the foot is not accelerating we get:
$F_{\text{normal}}=F_{\text{joint}}^{y}+m_{f}g$ (31)
This gives us an equation for $F_{\text{normal}}$:
$F_{\text{normal}}=ml_{b}(-\dot{\theta}^{2}s_{\theta}+\ddot{\theta}c_{\theta})+mg+m_{f}g$ (32)
Similarly, we can derive an expression for the ground reaction/friction force in the $x$ direction via force balancing:
$F_{\text{friction}}=F_{\text{joint}}^{x}=-ml_{b}(\dot{\theta}^{2}c_{\theta}+\ddot{\theta}s_{\theta})$ (33)
Finally, the location of the center of pressure (CoP) as a function of these known quantities can be found by balancing torques about the ankle:
$((l_{f}-a)-\text{CoP})\cdot F_{\text{normal}}-cm_{f}g+\tau-h_{f}F_{\text{friction}}=0$ (34)
which in turn gives us
$\text{CoP}=(l_{f}-a)-\frac{cm_{f}g-\tau+h_{f}F_{\text{friction}}}{F_{\text{normal}}}$ (35)
7.2 State Constraints
Constraints C3 and C5 can be converted into state constraints on $\tau$ by first solving 22 for the acceleration:
$\ddot{\theta}=\frac{1}{ml_{b}^{2}}\tau-\frac{g}{l_{b}}c_{\theta}$ (36)
[…]
$ml(-\dot{\theta}^{2}s_{\theta}+(\frac{1}{ml^{2}}\tau-\frac{g}{l}c_{\theta})c_{\theta})+mg+m_{f}g\geq 0.$ (37)
Solving for $\tau$, after some algebra we see that
$\tau\geq\frac{-l(m+m_{f})g+ml^{2}\dot{\theta}^{2}s_{\theta}}{c_{\theta}}+mglc_{\theta}$ (38)
if $c_{\theta}>0$, with the direction of the inequality flipped if $c_{\theta}<0$.
Similarly, plugging (36) into Constraint C5 results in four inequalities constraining $\tau$. Letting $\rho_{1}=\mu c_{\theta}+s_{\theta}$ and $\rho_{2}=\mu c_{\theta}+s_{\theta}$, when $\rho_{1}<0$
$\tau<mglc_{\theta}+\frac{(\mu s_{\theta}-c_{\theta})ml^{2}\dot{\theta}^{2}-\mu lg(m_{f}+m)}{\rho_{1}}$ (39)
with the direction of the inequality flipped when $\rho_{1}>0$. When $\rho_{2}<0$,
$\tau<mglc_{\theta}+\frac{(\mu s_{\theta}+c_{\theta})ml^{2}\dot{\theta}^{2}-\mu lg(m_{f}+m)}{\rho_{2}}$ (40)
Lastly, constraints on $\tau$ are derived from C4 by first incorporating the parameters of the foot:
$0\leq\text{CoP}\leq l_{f}$ (41)
Plugging in our expression for the CoP, equation (35), gives us two corresponding inequalities:
$\tau<aF_{\text{normal}}+h_{f}F_{\text{friction}}+cm_{f}g$ (42)
and
$\tau>-(l_{f}-a)F_{\text{normal}}+cm_{f}g+h_{f}F_{\text{friction}}$ (43)
8 Target Set Computation
8.1 Computing the CLF
we can write the following optimization problem, where both constraints come from Definition 4:
[…]
s.t. $P>0$
$(A+BK)^{T}P+P(A+BK)<0$
This problem is nonlinear in $K$ and $P$, but it can be converted to a convex SDP. First, we multiply the second constraint by $P^{-1}$ on both sides, so that it becomes
$P^{-1}((A+BK)^{T}+(A+BK))P^{-1}<0.$ (45)
Letting $Q=P^{-1}$ and $Y=KQ$, we can now write an SDP with linear constraints
[…]
s.t. $Y^{T}B^{T}+QA^{T}+BY<0.$
We now have a CLF $E(x)=x^{T}Px$, where $P=Q^{-1}$, and a controller $u=Kx$, where $K=YP$.
8.2 Computing maximally scaled ellipsoids
Let $(x-x_{\text{eq}})^{T}P(x-x_{\text{eq}})$ be the CLF corresponding to a given equilibrium point. Let $A_{\text{con}}x=b_{\text{con}}$ represent a set of $n$ linear constraints. We wish to find the largest $c$ such that the ellipsoid =$\mathcal{E}=\{x|(x-x_{\text{eq}})^{T}P(x-x_{\text{eq}})\leq c\}$ is fully contained within the constraints. Denote $a_{i}$, $b_{i}$ the $i$th row of $A_{\text{con}}$ and entry of $b_{\text{con}}$, respectively. In the case of a single constraint hyperplane $a_{i}^{T}x\leq b_{i}$, it is clear that the largest possible $c$ should produce an ellipsoid that just touches the constraint at one point. This can be cast as a simple convex optimization problem:
$\min_{x}\quad$ $(x-x_{\text{eq}})^{T}P(x-x_{\text{eq}})$ (47)
s.t. $a_{i}^{T}x=b_{i}$
By the method of Lagrange multipliers, we have that the Lagrangian $L(x)=(x-x_{\text{eq}})^{T}P(x-x_{\text{eq}})+\lambda(a_{i}^{T}x-b_{i})$. Taking the partial derivatives, we solve
$2P(x-x_{\text{eq}})+\lambda a_{i}^{T}$ $=0,$ (48)
$a_{i}^{T}x-b_{i}$ $=0$
for $\lambda,x$. After some algebra this gives an expression for $c$, which in turn should be minimized over all $n$ constraints:
$c=\min_{i}\,\frac{(b_{i}-a_{i}^{T}x_{\text{eq}})^{2}}{a_{i}^{T}P^{-1}a_{i}}$ (49)
9 Computational Considerations
To improve computational performance, the states of system (3) were nondimensionalized. Angular velocities $\dot{\theta}$ were scaled by $\omega=\sqrt{l/g}$, the natural frequency of the pendulum. The human torque state $\tau_{h}$ was nondimensionalized by multiplying by $\frac{\omega^{2}}{ml^{2}}$.
The backward reachable set was computed in the nondimensionalized state space over a $225\times 225\times 225$ grid. This was the coarsest grid which produced sufficiently clean results. We found that results computed over coarser grids were too noisy.

### v3

@@ 5.1 Stabilizable regions for age and sex adjusted models
Table 2 summarizes the results for the YF, YM, OM, OF, and wOF models with the GC and PD controller strategies. The changes in SR area for the forward and backward velocity regions are listed separately. In the forward angular velocity region, the GC and PD controllers slightly reduce the area of the SR in the young adult models but increase the total area in the OF and wOF model. The backward velocity region shows a different trend — the exoskeleton assistance reduces SR area for all models. The additional damping provided by the PD controller does not significantly mitigate reductions in SR area relative to gravity compensation alone. The stabilizable regions for the YM and the wOF models are shown in Figure 4a.
As seen clearly in the figure, the addition of exoskeleton torque does not uniformly increase or decrease the area of the SR. We see the following effects in the positive (posterior-to-anterior) velocity region:
  • The exo increases the SR along the entire backward fall boundary and along the forward fall boundary at low velocities.
@@
At negative (anterior-to-posterior) velocities we see:
  • The exo increases the SR along the forward fall boundary up to nondimensional velocities of approximately $-0.7$ and along the backward fall boundary at low velocities.
@@
Table 2: Percent change in total area of stabilizable region for sex- and age-adjusted models with exoskeleton assistance relative to no-exo baseline.
Gravity compensation Proportional-derivative
$v^{+}$ $v^{-}$ $v^{+}$ $v^{-}$
YM $-0.4$ $-16.9$ $-0.4$ $-16.5$
YF $-3.1$ $-32.7$ $-2.7$ $-32.4$
OM $+3.7$ $-15.4$ $+4.6$ $-14.9$
OF $+8.4$ $-19.6$ $+9.6$ $-18.9$
wOF $+15.9$ $-10.5$ $+17.5$ $-10.2$
5.2 Effect of reduced MT and MRTD
A graphical representation of the stabilizable region for YF, with and without exoskeleton assistance, when MT or MRTD are reduced by $30\%$ and $50\%$, is shown in Figures 4b and 4c. At positive CoM velocities, reducing maximum torque causes large reductions in SR area, but the addition of exoskeleton assistance mitigates the overall effect. For example, a $30\%$ MT reduction results in an SR area that is $20\%$ smaller than the baseline region. The SR area is only $10\%$ smaller when ankle exoskeleton assistance is included (Fig. 5a). At negative velocities, the exoskeleton exacerbates MT deficits (Fig. 5b).
When MRTD is reduced, the effect on the overall SR only becomes noticeable after large reductions ($\geq 40\%$) at positive velocities. When an exoskeleton is added, however, the effect is amplified. Without the exoskeleton a $30\%$ reduction in MRTD causes a $2.6\%$ reduction in SR area. With an exoskeleton the SR area is reduced by $10.4\%$. At negative velocities, the change in total area is dominated by the reduction induced by simply adding the ankle exoskeleton.
Figure 5: Trends in stabilizable region area when maximum torque (MT, circular markers) and maximum rate of torque development (MRTD, diamond markers) are independently reduced, with (filled/purple) and without the exoskeleton (open/orange). 5a) At forward velocities, the addition of the exoskeleton amplifies MRTD deficits, while it mitigates MT deficits, at least with respect to the % SR area measure. 5b) At backward velocities, the exoskeleton tightens the foot-ground interaction constraints. This results in a large reduction in % SR area, compared to the more subtle changes related to reduced MT or MRTD.
5.3 Work-energy bounds for the simpler model
Figure 6: Results of the work-energy (W-E) analysis. The zero torque line (ZTL) is shown in (6a) for the system without (orange line) and with the exoskeleton (purple line). Along these lines the human does not need to produce any torque to reach static equilibrium. Corresponding lines indicating the boundaries of the stabilizable region (SR), computed using W-E analysis, are shown in (6b). The SRs computed with our reachability-based method without (orange region) and with the exoskeleton (purple region) are shown for comparison. While the W-E analysis can capture some of the effects of ankle exoskeleton assistance, it cannot account for foot-ground interaction constraints or limited torque production capabilities and overapproximates the boundaries of the SR.
Near the upright posture and at low ankle angular velocities, the stabilizable regions for the exoskeleton and non-exoskeleton systems are well approximated by the bounds derived from the work-energy analysis (shown in Figure 6b for the wOF). At higher velocities and larger ankle angles, the work-energy bounds overapproximate the SR due to their inability to account for the foot-ground interaction constraints and the limited human MT and MRTD.
6 Discussion
In this study, we sought to characterize the effect of exoskeleton assistance at the ankle on feasible standing stability. We used simple, age- and sex-specific models, that reflect differences in human ankle torque production capabilities. The age related changes include reduced MT and MRTD. For each model, we computed the SR, consisting of all CoM positions and velocities from which it is feasible to recover to a quiet standing position without taking a step. We found that added ankle exoskeleton assistance with standard control strategies can increase feasible stability in low velocity nominal conditions for weaker models, but that it hinders stability for all models at higher CoM velocities.
The SRs computed without ankle exoskeleton assistance show that reduced MT has a far stronger effect on feasible stability than MRTD. Even a severe reduction in MRTD resulted in only a $10\%$ reduction in total SR area (Figure 5). This aligns with results from the literature showing that the difference in plantar and dorsiflexor MRTD between fallers and non-fallers is either weakly [13] or not at all [12] significant. Interestingly, the stabilizable regions computed with ankle exoskeleton assistance suggest that the device may exacerbate some deficits. Reductions in MRTD lead to greater reductions in the total SR area when ankle exoskeleton assistance is included as compared to the baseline results without the exoskeleton (diamond shaped markers in Figure 5a and 5b). For reduced MT, the same detrimental effect is found only at negative (anterior-to-posterior) velocities; at positive (posterior-to-anterior) velocities, the ankle exoskeleton mitigates the reduction in SR area.
In full strength models, it is perhaps not surprising that the addition of the exoskeleton does not increase the area of the SR. These models have sufficiently high MT and MRTD bounds to maximize the size of the SR within the region where the foot-ground interaction constraints are not violated. The lack of change in total SR area at forward velocities (Table 2) aligns with the results from [20], which showed no change in CoM kinematics as a result of moderate forward perturbations for young adults wearing ankle exoskeletons using the same controllers modeled here.
At negative velocities, the SR of the full strength human-exoskeleton models is shrunk primarily along the short edge of the nominal SR (lower right quadrant of the subplots in Fig. 4). This is due to the exoskeleton shifting the constraints on the total human torque (Equation (4)). It must hold that $h_{\text{lb}}(x)-\tau_{\text{exo}}<x_{3}<\tau^{\text{max df}}_{\text{human}}$ when the left hand side is positive, but maximum (positive) dorsiflexion torque magnitude can be as low as $25\%$ of maximum (negative) plantar flexion torque, even in full strength models. This can result in constraint violations when $\tau_{\text{exo}}$ is large and negative.
@@ 6 Discussion
In weaker models, the SR area measurement cannot fully describe the effects of an exoskeleton on standing balance as the exoskeleton does not uniformly alter the SR boundary (Fig. 4). As noted in section 4.4, gravity sometimes provides an assistive torque in the no exoskeleton case. Therefore, if the exoskeleton controller is designed to cancel the gravitational torque, we might expect that exoskeleton “assistance” does not uniformly improve stability. On the other hand, the exoskeleton increases the torque bounds of the human user. This is more important for weaker models, as stronger models can already produce enough torque to maximize feasible stability and are mainly limited by foot-ground interaction constraints. In particular, an exoskeleton increases static stability in weaker models, resulting in a larger functional base of support. A weak adult wearing the exoskeleton has the ability to do more mechanical work, but they must compensate at times for the gravitational torque being canceled out by the exoskeleton.
The bounds computed using work-energy analysis approximate when this trade-off, between the exoskeleton providing mechanical work vs. requiring users to do additional mechanical work, occurs. Thus, these bounds indicate regions in the phase plane where exoskeleton assistance drives the reduction in feasible stability (Figure (6)). Work-energy bounds are, however, far less precise than the stabilizable regions, because the work-energy analysis assumes conservation of energy, ignores ground contact and input constraints, and cannot generalize to systems with dissipation (such as exoskeleton controllers with damping). Together, the higher and lower dimensional models provide a unique way to understand the mechanisms through which ankle exoskeletons may affect standing balance. As we find a benefit to having an exoskeleton in low velocity conditions, but a corresponding cost at high velocities, a major implication of our work is that while designing exoskeleton controllers, we must consider what nominal conditions and extreme scenarios older users might experience and then test these devices accordingly. There may be unintended consequences from an assistive device that aids stability in most cases if it increases the risk of falls in even a small percentage of scenarios a user may experience.
A limitation of this study is the simplicity of the models. Our results generate bounds on ankle strategies for balance, but they do not allow the use of hip or knee strategies that could enhance the SR. In that sense, our results are likely an under-approximation of the true SR. We model both the human and exoskeleton torque simply, not accounting for nonlinear muscle dynamics, physiological actuation delays, device delays, and the physical interface between the human body and the device. The interaction between physiological delay and exoskeleton actuation delay, in particular, likely affects ankle exoskeleton-assisted standing balance [14, 15], so we plan to include delay in future modeling work.
An interesting result of the work-energy analysis relates to the shifts in the ZTL when the exoskeleton is added. Examining the ZTL with the exoskeleton (Fig. 6a, purple lines), we can see that in the forward velocity half of the phase plane the line is shifted to the left of the no-exoskeleton ZTL (Fig. 6a, orange line). Thus, a plantar flexion braking torque is always required to prevent forward falls. The addition of the exoskeleton effectively eliminates the possibility of a backward fall occurring, as long as there is an initial positive velocity and the ground contact constraints are satisfied. In the negative velocity half of the plane, the ZTL is shifted to the right, thereby increasing the area where a dorsiflexion braking torque is needed to prevent a backward fall. These shifts in the ZTL mean that the exoskeleton qualitatively changes the optimal human response required to stabilize to quiet standing in large regions of the state space. This may have unforeseen implications on the ability of users to adapt to exoskeleton assistance, preventing them from responding with optimal control strategies. We plan to explore this idea further by modeling suboptimal adaptation to these devices, which could help guide future empirical testing of device controllers.
It is also important to analyze more sophisticated controller strategies. The small reduction in area in the full-strength YM model results from having a large enough range of positive and negative MT and large enough MRTD such that undesireable exoskeleton behavior can be compensated for quickly. An exoskeleton that compensates for missing MT and MRTD by directly amplifying ankle torque could enable older users to uniformly enlarge their SR. On the other hand, such a strategy could unintentionally amplify suboptimal human control strategies. This approach is often implemented as proportional myoelectric control [36]. To enable analysis of such myoelectric controllers, we will include more detailed muscle actuation models in our future work.
Lastly, we hope to develop model-based experimental protocols. To fully understand the effect of ankle exoskeletons on stability, the stabilizable regions suggest that it is insufficient to apply relatively moderate forward perturbations that are common in experiments [20, 14]. Our regions can serve as a guide in the design of perturbation experiments of older adults wearing exoskeletons. Given the torque production capabilities of a participant, the perturbation magnitude and direction required to sufficiently challenge stability can be calculated. This will enable thorough, model-driven experimental evaluations of ankle-exoskeleton assisted standing balance.